# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ 3.1 Model Performance Overview
      Figure 5: F1 scores for high-level labels based on the best-performing configuration for each model. For the Migrant group, the two test sets (Test 1 and Test 2) are combined.
    • Performance drops on historical data. All models perform worse on the historical Migrant (T2) set, including the strongest ones. This reflects a broader pattern: model performance tends to decline on older data, as shown by macro F1 scores over time (see Figure 16)—so does agreement among human annotators (Cohen’s Kappa). However, while human trends are smooth and monotonic, model trajectories are less stable. Stronger models such as GPT-4, Mistral-2-Large, Llama-3.3-70B, and BERT show a general upward trend but exhibit clear dips in the 1930s and 1960s. In contrast, smaller models (e.g., Llama-3-8B, SBERT, Gemma-9B) show little to no temporal coherence.
In summary, GPT-4 remains the strongest model overall. While performance generally scales with model size, mid-sized models can approximate GPT-4 when given examples. Among open-weight alternatives, Llama-3.3-70B emerges as the most viable option, particularly for the Migrant category, delivering competitive results in both high-level and fine-grained classification, while offering a more cost-effective solution. We therefore adopt it for the subsequent social scientific analyses. However, since it still lags behind GPT-4 in capturing certain fine-grained subtypes, we further validate its reliability by comparing the temporal patterns produced by both models in 4.1.
3.2 Error Analysis
In this subsection, we examine the errors made by the two strongest models in our experiments, GPT-4 and Llama-3.3-70B, as determined by macro F1 scores. To identify where these models struggle, we compare their confusion matrices on the full Migrant test set (Figure 6 for high-level; see Figure 17 in the Appendix for confusion matrices for all models included in the study and Figure 18 in the Appendix for fine-grained label confusion matrices) with a human reference, constructed by averaging confusion matrices from four leave-one-out (LOO) comparisons among annotators.
Figure 6: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four LOO comparisons, where each annotator is compared to the consensus of the others.
Both models exhibit notable differences from human annotators. In particular, both models frequently annotate true Solidarity instances as None (108 instances for GPT-4, 128 for Llama) and true Anti-solidarity instances as None (59 for GPT-4, 60 for Llama), especially for common subtypes such as group-based and compassionate. These errors are far less frequent in human annotations (32 and 12, respectively), suggesting that models may more often underestimate both solidarity and anti-solidarity. Llama additionally shows a tendency to misclassify true Solidarity as Anti-solidarity (30 cases vs. 11 for GPT-4 and 13 for human annotators) and, to a lesser extent, true Anti-solidarity as Solidarity (18 vs. 4 for GPT-4 and 10 for humans). Some of these misclassifications arise when a speaker cites an anti-solidarity statement by another party, which Llama occasionally misinterprets as anti-solidarity (see Examples 9 and 11 in Table 14 in the Appendix; these are discussed in more detail in Section 4.3. However, manual inspection indicates that such cases are rare.
@@ 4 Temporal Trends in Solidarity Towards Migrants
In this section, we apply the best-performing open-weight model identified in 3, Llama-3.3-70B, to the full Migrant dataset of 63k instances to analyze trends across time. We focus on the Migrant dataset in particular because migration has long been one of the most contested issues in German politics, consistently framed through ?crisis narratives? that have shaped national identity and migration governance across decades frohlich2025migration (36). This makes it a compelling case for evaluating whether model-based annotations align with historical and political science literature. It is also reflected in our human-annotated dataset, where migrant-related instances display a more diverse label distribution, including substantially more anti-solidarity, compared to the predominantly solidarity-oriented discourse on women. More specifically, 4.1 compares the long-term (anti-)solidarity trends across human annotators and LLMs and thereby further validates the results of our models. In doing so, we also note two periods with notable shifts in solidarity, namely the post-war period and the recent past, which we cover in more detail in Section 4.2 and Section 4.3.
(a) GPT-4
(b) Llama-3.3-70B
Figure 7: Solidarity trends for Migrant predicted by GPT-4 and Llama-3.3-70B. Values are normalized within each decade; percentages may not total 100% because instances classified as ?None? are not shown (though included in the calculation). The period 1933–1949 (grey shaded area) is excluded due to limited data availability during and immediately after the Nazi dictatorship.
4.1 Trends with GPT-4 vs with Llama-3.3-70B
(a) GPT-4
(b) Llama-3.3-70B
Figure 8: (Anti-)solidarity frames trends for Migrant predicted by GPT-4 and Llama-3.3-70B. Percentages are normalized within each decade across all eight subtypes; thus, the curves in both panels together sum to 100% per decade. The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
Figures 7 and 8 show the development of (anti-)solidarity for GPT-4 and Llama-3.3-70B on a high and fine-grained level across 155 years. The GPT-4 reference data is taken from kostikova-fine (57), based on a subset of 18.3k migrant-related instances (from an original 58k-instance dataset; see 2.1) spanning 1867–2022. While Llama-3.3-70B is applied to the full dataset (expanded to 63k instances through 2025) for the main analysis, we restrict it to the same 18.3k-instance subset for this comparison.
To validate the annotations of the models, we first compare them to each other. Comparing the trends between GPT-4 and Llama-3.3-70B, we observe a high correlation of 0.96 on the high level and between 0.77 and 0.95 across the more frequent fine-grained subtypes (see Table 4 in the Appendix). This indicates that the models show a high degree of agreement in temporal trends. [^5] Additionally, Llama-3.3-70B shows near-identical trends between the 18.3k subset and the full 63k dataset (e.g., $\rho$ = 0.996 for solidarity and anti-solidarity, and $\sim$0.97–0.98 for subtypes).
For reference, Figure 3 shows the high-level trends across time for human annotations ($N=2,084$). We observe similar relative distributions of (anti-)solidarity on a high level. Direct trend correlations between Llama-3.3-70B and human annotations are also substantial ($\rho$ = 0.84 for anti-solidarity, $\rho$ = 0.76 for solidarity, $\rho$ = 0.47 for mixed). In terms of Cohen’s $\kappa$ (computed on the 2,084 human annotated instances), GPT-4 slightly outperforms Llama-3.3-70B ($\kappa$ = 0.57 vs. 0.50 at the high level; 0.38 vs. 0.36 for subtypes), but differences are modest. Overall, GPT-4 and Llama-3.3-70B appear to produce similar results that agree to a moderate level with human annotations. Given that it is an open weight model that we can run on our own computing hardware at low further cost, we utilize Llama-3.3-70B for the subsequent analysis.
4.1.1 Trend Stability Across Keywords.
Before we move on to interpreting the observed trends derived from Llama-3.3-70B predictions, we verify that they are not artifacts of our keyword-based data collection (see 2.1). As Figures 19 and 20 in the Appendix show, individual keywords vary considerably in their temporal distribution and in the balance of solidarity vs. anti-solidarity. To rule out the possibility that our long-term trends merely reflect such keyword effects, we perform two robustness tests.

### v2

@@ 3.1.2 Task and label-related challenges
    Figure 5: F1 scores for high-level labels based on the best-performing configuration for each model. For Migrant, the two test sets (Test 1 and Test 2) are combined.
  • Performance drops on historical data. All models, including the strongest, perform worse on the historical Migrant (T2) set. This reflects a broader trend: performance declines on older data (see Figure˜18 in the Appendix), as does human agreement (Cohen’s Kappa). We examine errors on the historical data in more detail in Section˜3.1.4.
3.1.3 Model Ensemble Performance
We construct a majority-vote ensemble from the best-performing open models in Table˜1: Llama-3.3-70B, Qwen-2.5, and gpt-oss-120B. We exclude GPT-5 due to its higher cost and because the subsequent social-scientific analysis focuses on open models.
We find that the ensemble does not outperform the strongest individual model (gpt-oss-120B), indicating that simple majority voting does not consistently improve performance in this setting (see Table˜2). This is in line with prior work showing that ensemble methods are most effective when individual models achieve comparable performance while making at least partly uncorrelated errors (Trad and Chehab, 2024).
Model Woman Migrant (T1) Migrant (T2)
gpt-oss-120B (Few-shot) 0.38 (0.55) 0.48 (0.69) 0.28 (0.57)
Llama-3.3-70B (Few-shot) 0.30 (0.50) 0.38 (0.65) 0.29 (0.58)
Qwen-2.5 (Few-shot) 0.22 (0.46) 0.43 (0.59) 0.27 (0.51)
Ensemble 0.34 (0.60) 0.47 (0.70) 0.29 (0.60)
Table 2: Performance (macro F1) of models included in the ensemble and the resulting majority-weighted ensemble on fine-grained and high-level tasks (in parentheses). In tie cases, votes are weighted by validation-set macro-F1. Tie-breaking weights are based on validation-set performance, ranking gpt-oss-120B >> Llama-3.3-70B >> Qwen-2.5 for the migrant setting and Llama-3.3-70B >> gpt-oss-120B >> Qwen-2.5 for the woman setting.
In summary, GPT-5 achieves the highest overall performance, surpassing human annotators on fine-grained level and approaching them on the high-level task. However, its closed-weight nature and high cost make it less suitable for our large-scale social scientific analysis. Among open models, gpt-oss-120B performs comparably to GPT-5 and approaches human-level performance (even matching it on the fine-grained level), making it the strongest open-weight alternative.
That said, the fact that these models reach or exceed human reference values does not imply that they can be used as direct substitutes of human annotators. First, the observed macro-F1 scores remain low enough that substantial and systematic model errors can still occur, including errors that human annotators are less likely to make; moreover, recent work shows that even high agreement with human consensus can coexist with systematic misclassification patterns (Bavaresco et al., 2025). We therefore examine these patterns through error analysis (Section˜3.1.4).
Second, in our setup, the human reference is derived from leave-one-out agreement and therefore inherits the uncertainty present in the human annotations. As a result, perfect F1 may itself be unattainable because some disagreement among human annotators likely reflects genuine label variation. In a subjective task like ours, this motivates moving beyond hard labels to consider label distributions instead (Section˜3.2.1). We address both issues next.
3.1.4 Error Analysis
We analyze errors of the strongest models through their confusion matrices on the full Migrant test set, comparing them with a human reference aggregated from leave-one-out (LOO) annotator comparisons. [^5] Because the human panel compares each annotator to a LOO consensus, whereas the model panels compare predictions to gold labels, we use row-normalized confusion matrices so that error patterns for each reference label are more directly comparable. Figure˜6 shows the top models; full matrices appear in Figure˜19 and Figure˜20 in the Appendix. We consider four aspects of model error: (i) high-level confusion; (ii) confusion within (anti-)solidarity subtypes; (iii) errors in historical instances (1867-1900); and (iv) ensemble failure analysis, with particular attention to error correlation across models. Since our substantive analysis focuses on migrant-related discourse, we restrict the error analysis to the Migrant category.
(a) High-level confusion matrices.
(b) Fine-grained confusion matrices.
Figure 6: Row-normalized confusion matrices (%) for the best-performing configuration of each of the top models by macro-F1, as reported in Table˜1, on the two Migrant test sets combined (Test 1 and Test 2). Each row sums to 100, showing how items with a given reference label are distributed across predicted labels. Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators. The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
(i) High-level confusion
Most models recover the broad high-level structure of the task. GPT-5 and gpt-oss-120B show the clearest separation between solidarity, anti-solidarity, and none, but even these stronger models do not fully match human error patterns. Relative to the human LOO panel, models systematically overassign cases to none, suggesting that they often fail to detect low-intensity or indirect stance expressions, despite the prompt instruction to consider even slight expressions of solidarity regardless of the main topic of the text Example 11 in Table˜14).
The models also preserve mixed much less often: for reference-mixed cases, the share predicted as mixed is 42% for humans, compared with 27% for GPT-5, 22% for gpt-oss-120B and Qwen, and 33% for the ensemble; Llama-3.3-70B is the only exception, reaching 47%. Instead, models resolve these cases into one-sided solidarity or anti-solidarity labels, suggesting that models overlook weaker cues of the opposing stance, as in Example 11.
A smaller but qualitatively important error type arises when speakers from party A quote or paraphrase anti-migrant rhetoric from party B to criticize party B and its stance. In such cases, models sometimes misclassify the speaker’s stance as anti-solidarity (e.g. Llama and Qwen in Examples 11 and 12).
(ii) Confusion within (anti-)solidarity subtypes
Within solidarity, and focusing on the better-represented labels (Table˜6), human disagreement is greatest between group-based and compassionate solidarity: in the human LOO matrix, 18% of group-based solidarity cases are labeled as compassionate solidarity by the held-out annotator. Models show less confusion in this direction (5-7%), but more often predict group-based solidarity for human-labeled compassionate solidarity instead (e.g. 32% for gpt-oss-120B). Humans also identify compassionate solidarity more consistently overall (66% vs. roughly 40% for models).
The two subtypes can co-occur in migration debates. Human annotators reflect this ambiguity, whereas models tend to shift such cases toward group-based solidarity, apparently relying more heavily on cues such as integration, rights, and legal inclusion (Examples 12, 12, and 13). However, rights language alone does not imply group-based solidarity. When rights claims are framed in terms of vulnerability, deprivation, or need, compassionate solidarity is the stronger reading, as it aims to place the other ?in an equal position as I am?. When such claims are framed in terms of shared membership or common rights and duties, group-based solidarity is more appropriate (Thijssen, 2012; Thijssen and Verheyen, 2022).
A related pattern is observed for other subtypes: confusion occurs when broader integration language associated with group-based solidarity overlaps with characteristics of other frames, as in Example 13. Since Thijssen and Verheyen (2022) describe the categories as conceptually close and the coding as interpretive, we treat such variation as meaningful uncertainty rather than noise. Within anti-solidarity, the sharpest subtype confusions appear in the rarer categories, especially empathic and compassionate anti-solidarity. These cases are often collapsed into group-based anti-solidarity, suggesting that models over-rely on broad cues of protecting the interests of the native in-group (see Example 14). Nevertheless, some models also show confusion among the better-represented anti-solidarity categories such as exchange-based anti-solidarity and group-based anti-solidarity (22% cases for Llama-3.3-70B, 21% for humans, see Example 13).
(a) High-level confusion matrices.
(b) Fine-grained confusion matrices.
Figure 7: Row-normalized high-level and fine-grained confusion matrices (%) for the best-performing configurations of the top models on the combined Migrant test sets, restricted to 1867-1900. As in Figure˜6, rows sum to 100, Human Annotators (LOO) denotes aggregated leave-one-out annotator comparisons, and the Ensemble aggregates Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
(iii) Confusion for historical instances
The historical sample (restricted to 1867-1900, Figure˜7) shows an even stronger overprediction of none. Across nearly all models, errors for many true classes collapse into none, more than in the human LOO panel; this is particularly visible for Qwen-2.5-72B, Llama-3.3-70B, and the ensemble. Human disagreement also increases on the historical subset, butannotators more often distribute errors across neighboring stance categories. Examples 14 and 14 indicate why this matters for historical inference. Here, solidarity is expressed through legal and constitutional reasoning rather than overt emotional language, and models fail to recognize such low-intensity stance at all.
Overall, strong models reproduce the broad label structure of the task, but their errors remain systematic rather than random. As a result, trends in (anti-)solidarity and mixed stance are likely to be underestimated, especially in older instances. Within subtypes, group-based solidarity is likely to absorb other solidarity claims, while rarer categories are often collapsed or disappear entirely. The ensemble mitigates these biases only modestly, if at all (e.g. it still overpredicts none and underpredicts mixed relative to humans). This highlights the need for correction techniques such as DSL, which we analyze in Section˜3.2.2.
(iv) Ensemble failure analysis
We begin by asking how often ensemble errors contain any useful signal at all. On the combined Test 1 and Test 2 data, the ensemble makes 1,112 errors. Of these, 747 (67.2%) occur on instances where all models are wrong, whereas only 365 (32.8%) occur when at least one model predicts the correct label. Thus, most ensemble failures occur on examples where none of the component models predicts the correct label.
The remaining 365 cases suggest that complementary signal exists in principle. To quantify how much performance could be recovered from this signal, we construct an oracle that selects a correct model prediction whenever at least one model is correct. The oracle reaches markedly higher macro-F1 than the ensemble–0.66 (0.87) on Test 1 and 0.45 (0.73) on Test 2, compared with 0.47 (0.70) and 0.27 (0.56), respectively. This indicates that although useful complementary signal is present in the model pool, simple ensembling fails to recover much of it in practice, which explains the ensemble’s substantially lower empirical F1.
A closer look at model overlap helps explain why this signal is difficult to exploit. Across the pooled Test 1 and Test 2 data, pairwise error Jaccard scores are consistently high (0.57-0.64; see Appendix˜D in the Appendix and Table˜4), indicating that models often fail on the same instances. Moreover, conditional on two models both being wrong, they predict the same wrong label in roughly half of cases (45-49%), which means that shared errors are often reinforced rather than corrected.
Taken together, these results suggest that the main limitation is limited error diversity across models on difficult instances. Section˜3.2 examines how the biases discussed in this section affect downstream trend estimates.
3.2 Improving Downstream Reliability
The preceding analyses identify two challenges for using LLM annotations in downstream inference. First, disagreement between human annotators does not necessarily reflect error, but may indicate intrinsic ambiguity and multiple defensible interpretations. Following work that discusses such cases under the umbrella term of label variation (Plank, 2022), we therefore move beyond hard labels and compare label distributions. Second, the error analysis shows that model predictions are systematically biased, with some labels over- or under-predicted relative to human annotations. Such bias can distort downstream analyses when predictions are aggregated into temporal or group-level trends. We address both issues in this section: we first assess the similarity of model and human label distributions (Section˜3.2.1), and then test whether debiasing with DSL improves recovery of human-derived trends (Section˜3.2.2).
3.2.1 Label distribution similarity
For each instance with at least three human annotations, we derive a human label distribution from the annotators’ votes. For the model ensemble, we analogously construct a label distribution from the votes of the three models (Llama-3.3-70B, Qwen-2.5, and gpt-oss-120B). For individual models, we use sharp distributions that assign 100% probability mass on the predicted label. We compare these distributions using Jensen–Shannon divergence (JSD) and total variation distance (TV), where lower values indicate higher similarity (see Appendix˜E in the Appendix for definitions). The analysis covers both the high- and fine-grained levels. We compare model–human distances against three baselines: (i) human-human distance, averaged over ten random annotator splits; (ii) a randomized model baseline, with predictions shuffled across instances; and (iii) single-model baselines comparing each model to the human distributions.
Table˜5 in the Appendix shows that the ensemble (TV 0.49; JSD 0.39) is closer to human label distributions (TV 0.45; JSD 0.35) than any individual model (TV 0.57–0.59; JSD 0.48–0.50) and far closer than the randomized baseline (TV 0.75; JSD 0.69), with the same pattern observed at the high level. These results indicate that strong LLMs reproduce human-like patterns of disagreement. At the same time, the ensemble–human distance does not surpass the human–human baseline.
These results have the following implications: First, since model disagreement parallels human uncertainty, this variation should be preserved through soft labels rather than reduced to single hard predictions. Second, because models do not surpass human–human similarity and their errors remain correlated, these soft labels still contain structured bias. As a result, model predictions cannot be used directly without correcting for bias. We address this issue in the next section.
3.2.2 Bias Correction with DSL
We validate all trend estimators considered below using the complete set of 2,184 Migrant-related instances that were annotated by humans (hereafter the evaluation subset; see Table˜6 in the Appendix). Trends computed from these annotations serve as the reference. Because the evaluation subset is fully human-annotated, we simulate a partial-label setting for DSL evaluation. Specifically, we randomly retain human labels for only a subset of instances within each decade and treat the remaining labels as unobserved. We then apply DSL using the corresponding decade-specific sampling probabilities and compare the resulting trends with those computed from the full human annotations.
We compare the human-derived reference trends against several alternative estimators. We first evaluate uncorrected estimators based on raw model outputs: (i) single-model hard-label predictions from the strongest models (Llama-3.3-70B, Llama-4-Scout, Qwen-2.5, gpt-oss-120B, GPT-5), (ii) a majority-vote ensemble of the best open models (Llama-3.3-70B, Qwen-2.5, gpt-oss-120B), and (iii) soft-label distributions obtained from these three open models’ outputs. We then evaluate debiased estimators obtained by applying DSL to (iv) a single strong model (Llama-3.3-70B) and (v) the soft-label ensemble. For each estimator, we compute temporal trends by grouping instances by decade and computing the share of instances assigned to each label in each decade. For DSL-based approaches, we use the bias-corrected pseudo-outcomes $\tilde{Y}_{i}$ directly in these aggregations. We then quantify agreement with the reference model by computing Pearson and Spearman correlations, as well as root mean squared error (RMSE), between model-based and human-based time series. We compute correlations separately for each label and report the average across labels.
Table˜3 shows that the soft-label ensemble corrected with DSL recovers human-derived temporal trends more accurately than other approaches, including a single-model DSL estimator, at both the high-level and fine-grained levels.
Model / Estimator Pearson Spearman RMSE
High-level trends
Llama-3.3-70B 0.87 0.86 0.08
Llama-4-Scout 0.74 0.76 0.08
Qwen-2.5 0.72 0.71 0.11
gpt-oss-120B 0.84 0.80 0.07
GPT-5 0.82 0.82 0.07
Majority vote ensemble 0.88 0.87 0.08
DSL (Llama-3.3-70B) 0.85 0.83 0.06
DSL (soft-label ensemble) 0.89 0.91 0.06
Fine-grained trends
Llama-3.3-70B 0.64 0.58 0.09
Llama-4-Scout 0.58 0.58 0.07
Qwen-2.5 0.36 0.38 0.11
gpt-oss-120B 0.44 0.52 0.10
GPT-5 0.56 0.50 0.08
Majority vote ensemble 0.64 0.67 0.10
DSL (Llama-3.3-70B) 0.50 0.49 0.07
DSL (soft-label ensemble) 0.68 0.67 0.04
Table 3: Comparison with human-based decade-level trends across estimators. The table reports average Pearson and Spearman correlations and RMSE (reference: human majority vote).
At the high level, the strongest single models and ensemble variants already achieve high agreement with human trends, with Pearson and Spearman correlations above 0.80 and low RMSE. In this setting, DSL shows moderate but consistent improvements: DSL applied to the soft-label ensemble achieves the highest correlations (Pearson 0.89; Spearman 0.91) and the lowest RMSE (0.06).
(a) Human majority vote
(b) gpt-oss-120B
(c) DSL (soft labels)
Figure 8: High-level (anti-)solidarity trends for the Migrant category under different labeling and correction strategies based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions (a) and (c) in Section˜3.2.2). The grey shaded area marks the period 1930–1949, excluded due to limited data availability during and immediately after the NS dictatorship. Figure˜23 in the Appendix shows the same trends for the remaining configurations.
In contrast, fine-grained trends are substantially harder to approximate from raw model predictions alone. Single-model and ensemble estimators show lower correlations and higher error: notably, even models that achieve the highest macro F1 scores in Section˜3.1.1 show only moderate agreement with human-derived fine-grained trends (for example, gpt-oss-120B reaches a Pearson correlation of 0.44 with an RMSE of 0.10, and the majority-vote ensemble reaches 0.51 with an RMSE of 0.08). This indicates that strong instance-level performance does not guarantee accurate recovery of fine-grained temporal trends. Here, DSL has a clear effect, particularly in reducing error: DSL applied to soft-label ensembles achieves the lowest RMSE (0.04) and higher correlations (Pearson 0.68) than all uncorrected approaches.
(a) Human majority vote
(b) gpt-oss-120B
(c) DSL (soft labels)
Figure 9: Fine-grained (anti)solidarity trends for the Migrant category under different labeling and correction strategies based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions (a) and (c) in Section˜3.2.2). The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship. Figure˜24 in the Appendix shows the same trends for the remaining configurations.
Figure˜8 and Figure˜9 illustrate these results. First, DSL reduces overuse of the None category and correspondingly increases estimated trends for solidarity and anti-solidarity, for example for Llama-3.3-70B. This pattern is also reflected in Table˜8 in the Appendix, which compares per-label correlations with human-annotated decade-level trends for Llama-3.3-70B and soft-label DSL. There, DSL reduces the RMSE for None from 0.13 to 0.07.
Box 3.1: Label Shares Over Time Each speech belongs to a time period $t$ (e.g. a decade or a year) with $N_{t}$ speeches and receives a high-level or fine-grained label $\ell$. The estimated label share $\hat{p}_{t,\ell}$ in period $t$ is defined as follows: (a) Hard-label trends (e.g. human majority vote, gpt-oss-120B): $\hat{p}_{t,\ell}=\frac{N_{t,\ell}}{N_{t}}$, where $N_{t,\ell}$ denotes the number of speeches in period $t$ assigned label $\ell$. (b) Soft-label trends [^6] Reported in Figure˜24 in the Appendix.: $\hat{p}^{\text{soft}}_{t,\ell}=\frac{1}{N_{t}}\sum_{i\in t}Y_{i,\ell}$, where $Y_{i,\ell}$ denotes the probability of label $\ell$ for speech $i$, obtained from the distribution of annotator or model labels. (c) DSL (soft-label) trends: $\hat{p}^{DSL}_{t,\ell}=\frac{1}{N_{t}}\sum_{i\in t}\tilde{Y}_{i,\ell}$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2.
At the fine-grained level, raw model outputs, including gpt-oss-120B, Llama-3.3-70B, Qwen-2.5, and their ensemble variants, show unstable trends, with dominance shifting between group-based, exchange-based, and compassionate categories across decades. DSL achieves more stable trends that follows with the human reference more closely, particularly within solidarity. Most notably, all raw model trends overemphasize group-based solidarity at the expense of compassionate solidarity. DSL corrects this imbalance and restores the dominance of compassionate solidarity observed in human annotations. Table˜8 in the Appendix shows that these corrections correspond to substantial error reductions, with RMSE for group-based solidarity decreasing from 0.20 to 0.06 and for compassionate solidarity from 0.20 to 0.08. A similar pattern is observed for fine-grained anti-solidarity. In raw model outputs, anti-solidarity is often limited to a single dominant subtype or disappears entirely; for example, empathic and compassionate anti-solidarity are absent from trends derived from Llama-3.3-70B predictions. DSL restores these subtypes and recovers differentiation between group-based and exchange-based anti-solidarity.
In sum, these results suggest that combining soft-label model outputs with DSL is a viable strategy for studying long-run trends when full human annotation is not feasible.
@@ 4 Temporal Trends in Solidarity Towards Migrants
In this section, we analyze temporal trends in the full Migrant dataset (63k instances) using DSL-corrected soft labels. We relate the resulting patterns to findings from political science, sociology, and history on migration to assess whether the temporal trends recovered by the DSL-corrected estimates are consistent with established interpretations.
For these analyses, DSL uses the decade-stratified human annotation sample of 2,184 Migrant instances (3.5% of the full corpus) to estimate bias-corrected label distributions for the full corpus of 63k speeches. Since human annotations were obtained under a decade-stratified sampling design (see Section˜2.2), we estimate DSL with decade covariates and decade-specific labeling probabilities $\pi_{i}$, and reuse the resulting pseudo-outcomes $\tilde{Y}_{i}$ for all analyses that condition only on time, including the sub-period analyses introduced in Section˜4.1.2. For analyses that condition on additional variables, such as party or keyword, we re-estimate DSL on the relevant analysis sample, since the prediction function $\hat{g}(Q_{i},X_{i})$ changes when additional covariates are included. In these analyses, party or keyword indicators enter the DSL prediction step as additional covariates alongside decade indicators, while labeling probabilities $\pi_{i}$ remain defined at the decade level, consistent with the annotation design.
We first present the DSL-corrected trends for the full dataset, summarizing the broad long-term pattern in parliamentary discourse about migrants and briefly assessing whether the aggregate trends are robust to variation in keyword composition. We then identify two periods with particularly notable shifts in solidarity–the post-war period and the recent past–which we analyze in more detail in Section˜4.2 and Section˜4.3.
4.1 Overall Solidarity Trends with DSL, 1867-2025
Figure 10: Solidarity trends for Migrant speeches computed using DSL-corrected soft labels (see definition (c) in Section˜3.2.2) on the full 63K dataset. Figure˜10(a) shows the high-level (anti-)solidarity trends, while Figure˜10(b) shows distribution of fine-grained (anti-)solidarity subtypes over time. The period 1930–1949 (grey area) is excluded due to limited data during and after the Nazi dictatorship.
The overall trends from 1867 to 2025 are shown in Figure˜10. Across most of the observable period, and especially from the beginning of the 20th century through the 2010s, solidarity exceeds anti-solidarity in parliamentary discourse about migrants. This pattern is driven mainly by compassionate solidarity, with group-based solidarity playing a less prominent but persistent role. Solidarity peaks in the postwar decades, while anti-solidarity remains comparatively low until the late twentieth century. From the 1980s onward, anti-solidarity rises (especially its group-based, compassionate, and exchange-based subtypes), with the sharpest jump occurring between the 2010s and 2020s.
Before turning to specific historical periods, we assess whether these aggregate trends are robust to changes in keyword composition over time.
4.1.1 Robustness to Keyword Composition
As Figure˜21 and Figure˜22 in the Appendix show, individual keywords vary widely in temporal distribution in the balance of solidarity vs. anti-solidarity associated with them. To ensure our long-term trends are not driven by such effects, we conduct two robustness tests.